OpenStudy (caishaax3):

Write the equation of the line that passes through (–1, 5) and has a slope of 3 in point-slope form.

OpenStudy (caishaax3):

A. y=3x+2 B. y=3x+8 C. y=-3x-2 D. y=-3x+8

OpenStudy (anonymous):

Point slope-formula: \[y-y_{1}=m(x-x _{1})\]

OpenStudy (anonymous):

(-1,5) is the\[\left( x _{1},y _{1} \right)\]

OpenStudy (anonymous):

And m=3

OpenStudy (anonymous):

Plug them in.

OpenStudy (anonymous):